OpenStudy (anonymous):

Is Kepler's constant the same for moons orbiting a planet as it is for planets orbiting the Sun? Why?

OpenStudy (vincent-lyon.fr):

What do YOU think?

OpenStudy (anonymous):

I think it is, because I'm fairly certain that the same formula is used in both cases. But I need more validation and explanation if could provide it?

OpenStudy (anonymous):

Mmmm I don't think so since the constant value depends on the mass of the central object..

OpenStudy (anonymous):

So what do you think it is?

OpenStudy (anonymous):

Hmmm?

OpenStudy (anonymous):

Do you think that the constant is the same for moons orbiting a planet as it is for planets orbiting the Sun? I think the answer is an obvious "yes", hence the term "constant"... but why?

OpenStudy (anonymous):

Okay so the law states: \[R ^{3}/T ^{2}\] This constant is for one system only, and each of them will have a distinct constant. So think about the gravitational force in our solar system: \[F = GMm/R ^{2}\] G is gravitational constant, M is the sun's mass, m is a planet's mass, and R is the distance between the two masses. \[GMm/R ^{2} = mv ^{2}/R\] m and R cancels out on both sides... \[GM/R = v\] v is the revolution velocity of the planet. The velocity of revolution is the circumference \[2 \pi R \]divided by T (period of revolution). \[GM/R = (2\pi R/T)^2 = 4\pi^2R^2/T^2 \] Rearranging the equation gives you \[R^3/T^2 = GM/(4\pi^2) \] On both sides, G, M, and pi is constant. Therefore, right hand side is constant, and the value of R^3/T^2 is constant for all planets in the solar system. However, M, the mass of the central object, will be different depending on the primary planet and such. That took a while. Hehe ^^

OpenStudy (anonymous):

Wow! Thanks for the response, lots of help